OpenStudy (anonymous):

can anybody help me with the steps to solve this? 9/10x-1=19/20x+4

OpenStudy (saifoo.khan):

\[\frac{9}{10x -1} = \frac{19}{20x + 4}\] Cross multiply, \[9(20x + 4) = 19(10x - 1)\] Can you solve further?

OpenStudy (anonymous):

not sure!! I assume just multiply in?

OpenStudy (anonymous):

(180x+4)=(190x-1)

OpenStudy (anonymous):

(180x+36)=(190x-19)????